Cos(5x)cos(2x) + sin(5x)sin(2x)


1
Expert's answer
2021-02-23T11:41:27-0500


Use the angle subtraction formula for cosine:

"\cos(\alpha-\beta)=\cos\alpha\cos\beta+\sin\alpha\sin\beta"

Let

"\alpha=5x", "\beta=2x".

Then

"\cos5x\cos2x+\sin5x\sin2x=\cos(5x-2x)=\cos3x"